# SAFES: Sequential Privacy and Fairness Enhancing Data Synthesis for Responsible AI

### v1

@@ 5.3 Results
Figure 3: Examples of the privacy (different lines) vs fairness (y-axis) vs utility (x-axis) trade-off in the Adult experiment. In each plot, each point on a line represents the mean and the error bar indicates ±$\pm 1$ SD over 35 repeats at a different privacy loss parameter $\varepsilon$ value $\in\{10^{-2}\;(\mbox{rightmost}),10^{-1.5},10^{-1},\ldots,10\;(\mbox{leftmost})\}$; the smaller $x$-axis values, the better the utility.
Figure 4: Examples of the privacy (different lines) vs fairness (y-axis) vs utility (x-axis) trade-off in the COMPAS experiment. In each plot, each point on a line represents the mean and the error bar indicates ±$\pm 1$ SD over 35 repeats at a different privacy loss parameter $\varepsilon$ value $\in\{10^{-2}\;(\mbox{rightmost}),10^{-1.5},10^{-1},\ldots,10\;(\mbox{leftmost})\}$; the smaller $x$-axis values, the better the utility.
The observations from Figures 3 and 4 on the privacy-fairness-utility trade-off can be summarized as follows. In addition, observations (1) to (3) below can also be seen in Supplementary Figures S1-S3 (Adult) and S8-S10 (COMPAS), while observations (4) and (5) can be seen in Supplementary Figures S4-S7 (Adult) and S11-S14 (COMPAS).
  1. (1)
    As expected, stringent privacy (small $\varepsilon$) and fairness (small $\eta$) parameters result in higher general utility loss (larger marginal TVDs and KS test statistics) and worse classifier prediction performance (accuracy, F1-score, AUC, etc.).
  2. (2)
    For small $\varepsilon$, DP noise is more influential on utility than the fairness parameter $\eta$; for larger $\varepsilon$, $\eta$ has a more noticeable impact, especially on the dataset utility. The fairness transformation generally has little impact on the classifier prediction performance, with or without DP guarantees.
  3. (3)
    Even for strong privacy guarantees and fairness constraints ($\varepsilon\in(0.05,1)$ and small $\eta$), SAFES synthetic data have similar utility to the original by several metrics, including the TVD of three-way marginals, which were not explicitly sanitized in the privacy step of SAFES.
  4. (4)
    In general, at very small $\varepsilon$, the SAFES synthetic data display more fairness than the original largely due to the large amount of noise injected for the DP guarantees but have poor utility. For $\varepsilon$ large enough to avoid this but still small enough to yield strong privacy guarantees, the synthetic data show significantly enhanced fairness for all three protected groups compared to the original that favors the privileged group, along with similar utility to the original.
  5. (5)
    For small $\eta$, fairness guaranteed by SAFES is robust to changes in privacy loss, allowing confident calibration of the privacy-utility balance while maintaining fairness. Additionally, the achieved fairness stabilizes as $\eta$ decreases.
Figures 3 and 4 suggest the FN rate balance in several cases is positive rather than negative. At first glance, this seems to indicate that the “unprivileged” sex group (female in Adult, male in COMPAS) is more likely to have the favorable outcome per this metric. However, for this metric, a positive value is rather an indication that more unprivileged than privileged individuals are incorrectly classified to the unfavorable outcome, which is still a disadvantageous result for the unprivileged group. Additionally, when measuring COD jointly on race+sex in the COMPAS experiment, the COD of the DP-only synthetic data (solid blue line) does not appear to converge to the COD of the original dataset (solid blue dot). This is an artifact of the AIM subroutine in SAFES when selecting from only one- and two-way marginals, as we did in our evaluations. Three-way relationships, if they exist, such as between race, sex, and recidivism, may not be fully captured (see, Figure 2 in Section “Details of study method and experiment procedures” for an example). We also note that the classifier utility (accuracy, F1-score, and AUC) at the baseline without privacy and fairness procedure are lower than the state-of-the-art for these datasets, though they outperform a trivial classifier for each dataset. This is due in part to the known scalability limitations (McKenna et al.,, 2022; Hu et al.,, 2024) for both AIM and TOT that limit considering too many attributes when synthesizing data, resulting in less predictors are available to fit classifiers than in other applications. That said, the performance of our classifier are similar to those obtained in other works on DP data synthesis using these datasets. For example, the $\approx 0.8$ accuracy in the Adult experiment (Supplementary Fig. S3) matches the accuracy obtained by Pujol et al., (2023); the $\approx 0.7$ AUC in the COMPAS experiment (Supplementary Fig. S10) outperforms the AUC by Pereira et al., (2024). Finally, TOT seems to increase the F1-score in the Adult experiment, which is counter-intuitive. We conjecture this was a spurious effect and false positive signal due to the number of metrics analyzed, especially considering that no other metric in either experiment yield similar observation. Nevertheless, it may be worth more in depth investigation in future work.
6 Discussion
@@
We have presented the SAFES procedure, which synthesizes data that simultaneously achieve DP guarantees and satisfy fairness constraints. We have performed experiments on two real datasets to evaluate the privacy, fairness, and utility trade-off in SAFES. The results clearly demonstrate that SAFES can be applied to real-world scenarios to synthesize and release datasets that satisfy both strong DP guarantees and improved downstream classifier fairness metrics, without significantly degrading the general utility of the dataset or limiting to specific downstream learning tasks. To implement SAFES in practice, we recommend choosing the smallest possible $\varepsilon$ that gives an acceptable utility metric for the target problem, then choosing the smallest possible $\eta$ that still permits a solvable optimization problem for the fairness transformation.
While SAFES is a general framework for all data types, we focused on the AIM DP synthesizer and the TOT fairness pre-processor that are limited to categorical and discretized numeric variables in the experiments. For both AIM and TOT, the number of marginals and marginal cell counts grow exponentially with the numbers of variables and categories in each variable; in addition, the number of constraints in the TOT optimization also increases with the data dimensionality. In future work, we will examine the applications of SAFES in datasets with mixed categorical and numeric variables, using more general statistical or deep generative models to generate synthetic data and developing new fairness data transformations permitting numeric data, though such attempts so far have not produced promising results (Pereira et al.,, 2024). We will also develop methods to improve the scalability of SAFES in high-dimensional settings. Another interesting direction for future research is to extend the SAFES framework to satisfy fairness constraints in settings other than binary classification, such as multi-class classification, regression, ranking, and clustering. One challenges in these extensions is a lack of a set of accepted meaningful and interpretable fairness criteria.
@@
The SAFES procedure has the potential for positive societal impacts in a wide range of fields (e.g., healthcare, hiring, criminal justice), where privacy and fairness considerations are necessary in the deployment of responsible AI. Organizations might consider integrating SAFES into their data-driven decision making pipeline, confident that they are gaining useful insights from synthetic data with guaranteed privacy and improved fairness. On a cautionary note, SAFES should not be used as a black box. It is important to understand the privacy and fairness requirement for a given problem before implementing SAFES; a lack of understanding of these implications can easily result in unsubstantiated claims about privacy and/or fairness, which would likely exacerbate the issue.
Data and code
The Adult (Becker and Kohavi,, 1996) and COMPAS (ProPublica,, 2016) datasets used in the experiments in this paper are publicly available on the UCI Machine Learning Repository and Kaggle, respectively. The code for the experiments in this paper can be found at https://github.com/sgiddens/SAFES.
@@ Acknowledgments
This work is supported by the University of Notre Dame Schmitt Fellowship and Lucy Graduate Scholarship to S.G. A portion of the work was carried out while S.G. was an intern at SandboxAQ, and S.G. acknowledges helpful discussion from supervisors Nicolas Gama and Sandra Guasch.
Supplementary Materials
Additional figures for the Adult and COMPAS experiments
We present several additional figures for both the Adult and COMPAS experiments. Supplementary Fig. 5 through Supplementary Fig. 11 are for the Adult experiment, while Supplementary Fig. 12 through Supplementary Fig. 18 are for the COMPAS experiment.
We note that the error bands for different values of $\eta$ in the fairness-only simulations can sometimes be fairly large, especially for the larger values of $\eta$ in our experiments. Larger $\eta$ mean more flexibility in the permitted dataset distortions, and hence greater variability in the learned randomized fairness transformation. Smaller $\eta$, on the other hand, would only permit very specific changes and thus have less variability. This is most noticeable in the $\eta=0.1$ case of the right-hand plot of Supplementary Fig. 6. However, in that case, the entire error band of p-values falls in the “fail-to-reject” region of the plot, so they all essentially provide the same information.
@@
Figure 7: Mean $\pm$ 1 SD (error bars and shaded regions) prediction performance of the logistic regression model trained on SPaFAS synthetic data for the Adult experiment.
@@
Figure 14: Mean $\pm$ 1 SD (error bars and shaded regions) prediction performance of the logistic regression model trained on SPaFAS synthetic data for the COMPAS experiment.
@@
Tabular results for the Adult and COMPAS experiments
We present the full set of tabular results for all privacy and utility metrics for both datasets. Supplementary Table 4 shows the results of the Adult experiment for various $\varepsilon$ privacy parameters with no fairness transformation. Supplementary Table 5 and Supplementary Table 6 show the results of the Adult experiment for various $\varepsilon$ privacy parameters with a fairness transformation using $\eta=0.1$ and $\eta=0.025$, respectively. Supplementary Table 7 shows the results of the COMPAS experiment for various $\varepsilon$ privacy parameters with no fairness transformation. Supplementary Table 8 and Supplementary Table 9 show the results of the COMPAS experiment for various $\varepsilon$ privacy parameters with a fairness transformation using $\eta=0.15$ and $\eta=0.08$, respectively.
$\varepsilon$ $10^{-2.0}$ $10^{-1.5}$ $10^{-1.0}$ $10^{-0.5}$ $10^{0.0}$ $10^{0.5}$ $10^{1.0}$ None

### v3

@@ Results
We note that the classification utility (accuracy, F1 score, and AUC) with the original data are lower than the state-of-the-art for these datasets, though they outperform a trivial classifier. This is partly because the original data includes only a subset of attributes and the continuous attributes were discretized to provide a fair baseline for assessing the utility loss from DP and fairness data preprocessings via AIM and TOT, both of which have known scalability limitations in terms of both number of attributes and the levels within an attribute and only work with categorical variables (thus the data preprocessing steps described in Section 4.1, resulting in fewer and less refined predictors being available to fit classifiers than would be possible if AIM and TOT were not applied.
4.3 SAFES (DP-CTGAN + RW) for data with mixed continuous and categorical data
We use the Adult data in Table 1 with continuous age and the German Credit data 2 for this experiment.
Privacy settings
We use the standard GAN objective with binary cross-entropy for both the critic and the generator without the gradient penalty term in the DP-CTGAN algorithm. The moment account method Abadi et al., 2016b is used as the privacy accountant $\mathcal{A}$ in Algorithm 3. We set the clipping bound $b=10$, batch size $m=500$, critic steps $s=1$, Adam learning rate $\alpha=2\times 10^{-4}$. For the Adult data, we examine $\varepsilon\in\{10^{-0.5},10^{0},\ldots,10^{2}\}$ and $\delta=10^{-7}$; the noise multiplier $\sigma$ corresponding to different $\varepsilon$ values are as follows: $\sigma=25$ for $\varepsilon\leq 0.5$; $\sigma=15$ for $0.5<\varepsilon\leq 1.0$; $\sigma=3.5$ for $1.0<\varepsilon\leq 3.5$; $\sigma=2.5$ for $3.5<\varepsilon\leq 10$; $\sigma=0.9$ for $10<\varepsilon\leq 50$; and $\sigma=0.5$ for $\varepsilon>50$. For the German Credit data, we examine one privacy loss at $\varepsilon=5$ and $\delta=5\times 10^{-5}$, the corresponding noise multiplier of which is $\sigma=3$.
Fairness settings
Unlike TOT, RW has no hyperparameters. Instead, it reweighs the training data distribution through loss function formulation by attaching an instance weight $W(G,Y)$ in equation (14) calculated from the data.
Experiment settings
In both the Adult (with continuous age) and German credit experiments, we split the original data (after the preprocessing steps described in Section 4.1) into a training set and a test set in an 75/25 ratio. Synthetic data of the same sample size as the training set are generated via SAFES in Algorithm 1 using DP-CTGAN as the DP data synthesizer; RW was applied for fairness when training the downstream classifier. We run 35 repeats on the Adult data and 15 repeats on German credit data, and report the average performance and stability across the runs. We use the SmartNoise SDK (OpenDP,, 2023) for the DP-CTGAN subroutine in the experiments, and the AIF360 library (Bellamy et al.,, 2018) for the RW subroutine. We train a logistic classifier on the DP synthetic data with RW and evaluate the prediction performance on the test set using the same set of metrics as in Section 4.2. For fairness evaluation, no COD is calculated as RW does not change the data; we calculate SPD, AOC, and the FP/FN rate balance for the trained logistic classifier with RW, where the protected attributes are race and sex separately and jointly in the Adult data and the protected attribute is sex in the German Credit data.
Results
A subset of the results for the Adult data are shown in Figs. 5 and 6; the rest can be found in the Appendix. Since RW does not alter the data, thus does not influence data-level metrics like TVD, Fig. 6 visualizes only the privacy-utility (i.e., ignoring fairness) trade-off for the general utility metrics. As expected, more stringent privacy decreases utility. Fig. 5 also shows that RW consistently improves fairness of the downstream classifier. The results for the German Credit experiment are presented in Table 4. First, RW for fairness barely affects the ML utility (Accuracy and F1 score) of the data but notably improves the fairness metrics (SPD and AOD) (the 2nd compared with the 1st rows in Table 4). Second, synthetic data via CTGAN, with and without DP, further improve the fairness on top of RW (the 3rd, 4th rows compared with 2nd rows). Third, DP reduces the ML utility of the synthetic data by 5% to 7% (the 4th vs. 3rd rows), but the decrease in the ML utility in the synthetic data generated by CTGAN – without DP (the 3rd vs. 2nd rows) is much more pronounced, representing a major challenge for GAN-based DP synthesizers in general. This is also reflected in Figs. 5 and 6, as well as noted in previous work (Pereira et al.,, 2024). That is, it is challenging to achieve near baseline utility with reasonable DP guarantees for GAN-based methods – even the largest privacy budgets in our experiments leave a noticeable utility gap compared to the baseline results. Lastly, in both the Adult and German Credit experiments, improved fairness due to DP noise (at the cost of low utility) is also present, as it was in the AIM + TOT experiments.
@@
Figure 5: Examples of the privacy (points on each line) vs fairness (y-axis) vs utility (x-axis) trade-off in the DP-CTGAN + RW experiment on the Adult data (with continuous age). In each plot, each point on a line represents the mean and the error bar indicates $\pm 1$ SD over 35 repeats at a different privacy loss parameter $\varepsilon$ value $\in\{10^{-0.5}\;(\mbox{rightmost}),10^{0},10^{0.5},\ldots,10^{2}\;(\mbox{leftmost})\}$; the orange dashed and blue solid lines correspond to with and without RW, respectively; for consistency with Figs. 2 and 3, $x$-axis values further left correspond to better utility.
Figure 6: Mean $\pm$ 1 SD (error bars) 3-way TVD and KS test statistic comparing original and synthetic datasets generated by SAFE(DP-CTGAN + RW) in the Adult data with continuous age.
Table 4: Mean (SD) ML utility and fairness metrics over 15 repeats for logistic regression classifier trained on synthetic German Credit data with and without DP and fairness RW.
Synthesizer Accuracy F1 score SPD AOD
Original 0.764 (0) 0.837 (0) -0.187 (0) -0.146 (0)
RW only 0.760 (0) 0.833 (0) -0.127 (0) -0.087 (0)
CTGAN (no DP)+RW 0.613 (0.085) 0.728 (0.104) 0.033 (0.067) 0.026 (0.064)
DP-CTGAN+RW (SAFES) 0.572 (0.117) 0.651 (0.186) 0.020 (0.068) 0.018 (0.057)
Overall, although DP-CTGAN permits non-categorical variables and does not suffer from the same scaling issues as the AIM + TOT approach for larger numbers of features, the utility of the synthetic data suffer, consistent with previous findings on GAN-based DP synthesizers. It is important to keep this in mind when using SAFES or a DP data synthesizer on data with continuous variables, as discretizing the variables may offer better privacy- utility tradeoff than adopting a GAN-based DP synthesizer.
4.4 Sensitivity analysis
Each algorithm – whether privacy-preserving or fairness-oriented – with the exception of the fairness RW method, involves hyperparameters. In particular, the TOT procedure has several domain hyperparameters, and DP-CTGAN also includes procedural hyperparameters (e.g., learning rate, batch size) that are generic to optimization and not specific to the privacy–fairness–utility trade-off emphasized in this paper. The experimental results in Sections 4.2 and 4.3 illustrate how the privacy-loss parameter $\varepsilon$ and the TOT fairness parameter $\eta$ (Section 4.2) shape the privacy–fairness–utility trade-off. Additional TOT domain hyperparameters that may affect this trade-off include the distortion function $\phi$ and the bounds on the probabilities of allowed transformation. Given $\phi$ is a dimensionless function with infinitely many possible specifications, we focus on analyzing the sensitivity of the fairness-utility trade-off with respect to $\mathbf{c}$ below.
Specifically, we fix privacy loss $\varepsilon$ at 1, keep the distortion function the same as in Section 4.2, and generate 35 datasets via SAFES (AIM + TOT) for each combination of $c_{1}\in\{0.2,0.15,0.1\},c_{2}\in\{0.1,0.075,0.05\},c_{3}\in\{0.05,0.025,0\}$, where $(c_{1},c_{2},c_{3})=\mathbf{c}$, and $\eta\in\{0.1,0.025\}$, leading to a total of $3\times 3\times 3\times 2=54$ scenarios. The results of the sensitivity analysis for a subset of utility metrics are presented in Fig. 7 (other general utility metrics, and performance metrics of a classifier trained on the synthetic data exhibited similar trends).
(a) F1 score Refer to caption
(b) 1-way TVD Refer to caption
(c) 3-way TVD Refer to caption
Figure 7: Mean $\pm$ SD (error bars) of F1 score (ML utility) and TVD (general utility) over 35 synthetic datasets generated via SAFES (AIM + TOT) at various values of $\mathbf{c}=(c_{1},c_{2},c_{3})$ in equation (10) in the Adult dataset.
The utility at $\eta=0.1$ – whether that’s general utility measured by TVD or the ML utility measured by the F1 score of the downstream classification – show robustness to varying $\mathbf{c}$ value. For the more stringent fairness setting of $\eta=0.025$, we observe a few trends. First, increasing $c_{1}$ (the probability of adjusting income from $\mathdollar 50k$ to $\geq$\$50k) has negligible impact on the ML utility of the downstream classifier, as seen in Fig. 7(a). For the general utility metrics measured by TVD, changing $\mathbf{c}$ affects the 1-way TVD in Fig. 7(a), but not the 3-way TVD presented in Fig. 7(b). In the case of 1-way TVD, increasing $\mathbf{c}$ (the probability of allowing small or large changes) in general decreases the utility, as expected, except a surprising observation that increasing $c_{1}$ (a greater probability of small changes) improves the 1-way TVD. We attribute this to TOT’s ability to optimize fairness by favoring numerous small adjustments over fewer large ones, thereby reducing the need for substantial transformations that may more negatively affect the general utility of the data. This conjecture is confirmed by the lower sensitivity observed in the 3-way TVD that jointly measuring across distributional differences over several variables in Fig. 7(b). This finding suggests a potential pathway to robust utility in TOT-transformed data, achieved alongside improved fairness (see Fig. 8, which shows that increasing $c$ in general improves fairness measured by SPD, as expected, especially at the more stringent fairness requirement of $\eta=0.025$.
Figure 8: Mean $\pm$ SD (error bars) SPD (fairness) over 35 synthetic datasets generated via SAFES (AIM + TOT) at various values of $\mathbf{c}=(c_{1},c_{2},c_{3})$ in equation (10) in the Adult dataset.
4.5 Runtime for SAFES (AIM + TOT)
As stated in previous sections, the main reasons for us to work with a subset of attributes from the Adult and COMPAS datasets for the AIM + TOT experiments in Section 4.2 is that both AIM and TOT face computational bottlenecks as the number of attributes increases. To further explore and quantify the computational scalability of SAFES (AIM + TOT) as implemented in our experiments, we evaluate the average runtime required to generate a synthetic dataset of different sample sizes and numbers of features, as displayed in Table 5. We observe that the runtime grows slowly with the sample size, but grows exponentially with the number of variables. This is a known limitation of both the AIM DP synthesizer and the TOT fairness data processor. For AIM, the number of marginals and marginal cell counts grow exponentially with the number of variables and categories of each variable. For TOT, the number of constraints in the optimization increases with the data dimensionality.
Table 5: Mean (SD) run time (seconds) over 30 runs of SAFES (AIM + TOT) on the Adult data of various sample sizes and attribute sets.
Features
{education, race, {education, race, {education, race,
Sample size income} sex, income} age, sex, income}
10000 6.748 (0.821) 11.493 (0.864) 21.032 (1.869)
20000 6.581 (0.398) 12.512 (1.199) 22.255 (1.488)
30000 6.911 (0.487) 12.618 (1.231) 25.617 (1.809)
40000 6.766 (0.272) 12.427 (0.783) 29.450 (3.919)
48842 7.163 (0.524) 12.459 (0.823) 34.591 (4.307)
5 Conclusions and discussion
@@
We have presented the SAFES procedure, which synthesizes data that simultaneously achieve DP guarantees and satisfy fairness constraints. We have run extensive experiments on three real datasets to evaluate the privacy, fairness, and utility trade-off in SAFES. The results demonstrate that in certain contexts, synthetic data generated by SAFES can enjoy strong DP guarantees and improved fairness for downstream classifier without significantly degrading the general utility of the dataset or constraining specific downstream learning tasks.
SAFES is a general framework. We demonstrated its usage with the AIM and DP-CTGAN synthesizers and the TOT and reweighting fairness data processor. Users of SAFES can employ any suitable DP synthesizer and fairness preprocessing method. Compared with the wide range of available DP synthesizers, fairness preprocessing methods are far less numerous. From a probabilistic ML perspective, we regard TOT as a principled approach despite its limitation in computational scalability; other data preprocessing methods for fairness tend to be more ad hoc and lack a strong probabilistic foundation. A comparison of different DP synthesizers and fairness data preprocessing methods is beyond the scope of this work but would be an interesting topic for future research.
For practical implementation of SAFES, we provide the following guidelines for choosing $\varepsilon$ for a DP synthesizer and $\eta$ in the TOT procedure. We recommend fixing the privacy parameter $\varepsilon$ first according to the desired privacy requirements, prevailing expectations, or common practice in the DP community, prior to preprocessing any sensitive dataset. For example, although $\varepsilon\leq 1$ is generally considered strong privacy guarantees, larger values of $\varepsilon$ have been used by government agencies or tech company to balance privacy and utility (e.g., $\varepsilon=19.61$ by the US Census Bureau (US Census Bureau,, 2021), $\varepsilon=2,8$ by Apple (Apple,, 2017), $\varepsilon=13.69$ by Google (Xu and Zhang,, 2024), etc. (Desfontain,, 2024)). For $\eta$ in the TOT procedure, our experiments indicate that the loss in utility as $\eta$ decreases is relatively small. A practical strategy is therefore to choose choose the smallest $\eta$ that still permits a feasible optimization (i.e., the problem remains solvable). That said, some applications may experience significant utility loss at small $\eta$, making the fairness–utility trade-off problem-specific. In such cases, the DP synthetic dataset produced by SAFES’s DP synthesizer can be used to test a range of $\eta$ values and identify an appropriate trade-off at no additional privacy loss.
In future work, we hope to develop methods to improve the scalability of SAFES to more complex data settings and diverse data types. Our experiment on the sensitivity analysis suggests that certain values of $\mathbf{c}$ in the TOT procedure seem to simultaneously improve fairness and utility. A more systematic investigation of this trend would be a promising direction. While our study focuses on binary classification – the predominant setting in the fairness literature and metrics Mehrabi et al., (2021), another interesting direction is to extend SAFES to satisfy fairness constraints in other settings, such as multi-class classification, regression, ranking, and clustering. A central challenge for these extensions, like in many fairness settings, is a lack of a set of accepted meaningful and interpretable fairness criteria.
@@
The SAFES procedure has the potential for positive societal impacts in a wide range of fields (e.g., healthcare, hiring, criminal justice), where privacy and fairness considerations are necessary in the deployment of responsible AI. Organizations might consider integrating SAFES into their data-driven decision making pipeline with confidence that they are gaining useful insights from synthetic data with guaranteed privacy and improved fairness. On a cautionary note, SAFES should not be treated as a black box. It is important to understand the privacy and fairness requirement of a given problem before implementing SAFES; a lack of understanding of these aspects can easily result in unsubstantiated claims about privacy and/or fairness, potentially exacerbating the very issues SAFES is designed to mitigate.
Code
The code for the experiments in this paper can be found at https://github.com/sgiddens/SAFES.
@@ Acknowledgments
This work was partially supported by the University of Notre Dame Schmitt Fellowship and Lucy Graduate Scholarship to S.G. A portion of the work was carried out while S.G. was an intern at SandboxAQ, and S.G. acknowledges helpful discussion from supervisors Nicolas Gama and Sandra Guasch.
Appendix
Appendix A Additional figures in the experiments
We present several additional figures for the AIM + TOT experiments with the Adult (Figure 1 through Figure 7) and COMPAS (Figure 8 through Figure 14) datasets, as well as the DP-CTGAN + RW experiment on the Adult (with continuous age) dataset (Figure 15 through Figure 21).
We note that the error bands for different values of $\eta$ in the fairness-only simulations can sometimes be fairly large, especially for the larger values of $\eta$ in our experiments. Larger $\eta$ mean more flexibility in the permitted dataset distortions, and hence greater variability in the learned randomized fairness transformation. Smaller $\eta$, on the other hand, would only permit very specific changes and thus have less variability. This is most noticeable in the $\eta=0.1$ case of the right-hand plot of Figure 2. However, in that case, the entire error band of p-values falls in the “fail-to-reject” region of the plot, so they all essentially provide the same information.
@@ Appendix A Additional figures in the experiments
Figure 3: Mean $\pm$ 1 SD (error bars and shaded regions) prediction performance of the logistic regression model trained on SAFES synthetic data for the Adult experiment.
@@
Figure 10: Mean $\pm$ 1 SD (error bars and shaded regions) prediction performance of the logistic regression model trained on SAFES synthetic data for the COMPAS experiment.
@@
Figure 15: Mean $\pm$ 1 SD (error bars and shaded regions) summed TVD in each marginal set for 1-way, 2-way, and 3-way marginals between the synthetic data vs the original data for the DP-CTGAN + RW experiment in the Adult data with continuous age.
Figure 16: Mean $\pm$ 1 SD (error bars and shaded regions) test statistic and corresponding p-value for the KS test comparing original and synthetic datasets for the DP-CTGAN + RW experiment in the Adult data with continuous age. Statistical significance threshold of $\alpha=0.05$ is marked in red in the plot on the right.
Figure 17: Mean $\pm$ 1 SD (error bars and shaded regions) prediction performance of the logistic regression model trained on SAFES synthetic data for the DP-CTGAN + RW experiment in the Adult data with continuous age.
Figure 18: Mean $\pm$ 1 SD (error bars and shaded regions) COD, measured with race, sex, and race+sex as the protected attribute, for the DP-CTGAN + RW experiment in the Adult data with continuous age.
Figure 19: Mean $\pm$ 1 SD (error bars and shaded regions) SPD, with race, sex, and race+sex as the protected attribute, for the DP-CTGAN + RW experiment in the Adult data with continuous age.
Figure 20: Mean $\pm$ 1 SD (error bars and shaded regions) FN rate balance, with race, sex, and race+sex as the protected attribute, for the DP-CTGAN + RW experiment.
Figure 21: Mean $\pm$ 1 SD (error bars and shaded regions) AOD, with race, sex, and race+sex as the protected attribute, for the DP-CTGAN + RW experiment.
Appendix B Tabular results
We present the full set of tabular results for all privacy and utility metrics for all experiments. Table 1 shows the results of the AIM + TOT experiment on the Adult dataset for various $\varepsilon$ privacy parameters with no fairness transformation. Tables 2 and 3 show the results of the AIM + TOT experiment on the Adult dataset for various $\varepsilon$ privacy parameters with a fairness transformation using $\eta=0.1$ and $\eta=0.025$, respectively. Table 4 shows the results of the AIM + TOT experiment on the COMPAS dataset for various $\varepsilon$ privacy parameters with no fairness transformation. Tables 5 and 6 show the results of the AIM + TOT experiment on the COMPAS dataset for various $\varepsilon$ privacy parameters with a fairness transformation using $\eta=0.15$ and $\eta=0.08$, respectively. Tables 7 and 8 contain the results of the DP-DP-CTGAN + RW experiment on the Adult dataset without and with the reweighting fairness data preprocessing, respectively.
Table 1: Full set of simulation results for the Adult dataset for various values of the privacy parameter $\varepsilon$ without the fairness transformation. The other privacy parameter $\delta$ is assumed to be $\delta=10^{-9}$ throughout. Values in the table are in the format “mean (SD)” of the output values from many runs of the evaluations. At least 30 runs were performed in each case. The final column represents a run that is also without the DP synthesis step as a baseline.
$\varepsilon$ $10^{-2.0}$ $^{-1.5} 101.0$$^{-1.0}$ $10^{-0.5}$ $10^{0.0}$ $^{0.5} 101.0$$^{1.0}$ None